a line through (-9,-9) and a slope of -7 has a y-intercept at what?
Answer by algebrapro18(206)

Their are 2 ways to solve this problem.

Method 1. Using the slope-intercept formula of a line.

We know that the equation of a line in slope interecept form is y=mx+b where m is the slope and b is the y intercept. Since we have the slope and a point we could plug those in and solve for b.

y = mx+b
-9 = 7(-9)+b
-9 = -63+b
54 = b

So we know the y-intercept is at 54.

Method 2. Using the point slope formula.

There is another equation that will give us the equation of a line. Its y-y1=m(x-x1). Where (x1,y1) is the given point and m is your slope.

Plugging in the values we have we get:

y-(-9) = 7(x-(-9))
y + 9 = 7(x+9)
y = 7x+63-9
y = 7x+54

SO we know that the y-intercept is 54.


Either method you use will get you the right answer.